OpenStudy (anonymous):

find the derivative of y with respect to x y=ln((1+sqrtx)/(x^5))

OpenStudy (lgbasallote):

quotient rule

OpenStudy (mimi_x3):

\[\ln\left(\frac{1+\sqrt{x}}{x^{5}}\right) \]?

OpenStudy (paxpolaris):

quotient rule ..... after chain rule

OpenStudy (anonymous):

yes mimi_x3

OpenStudy (turingtest):

I would avoid the quotient rule by rewriting the argument of the log

OpenStudy (paxpolaris):

nice

OpenStudy (turingtest):

\[\ln(x^{-5}+x^{-9/5})\]who need to make life harder? ;)

OpenStudy (paxpolaris):

i thought you meant: \[y=\ln \left( 1+\sqrt x \over x^5 \right)=\Large \ln \left( 1+\sqrt x \right)-5\ln \left( x \right)\]

OpenStudy (turingtest):

oh that's even better :)

OpenStudy (anonymous):

The way paxpolaris did it was an easy solve but it does not lead to one of the answers in the multiple choices.

OpenStudy (paxpolaris):

rationaliz denominator .... and maybe you need to combine the 2 fractions

OpenStudy (paxpolaris):

\[\large y\ '={1 \over 1+ \sqrt x}\times \frac 12 \cdot \frac 1 {\sqrt x}-\frac 5x\]

OpenStudy (anonymous):

what i ended up with was \[y=(1/(2x ^{1/2}(1+\sqrt{x})))-(5/x)\] if that is right... but its not an option on the homework choices.

OpenStudy (paxpolaris):

yes that's right ... is the denominator in at least one of you choices 2x(x-1)

OpenStudy (paxpolaris):

\[\implies \Large y\ ' ={1 \over 2 \left(x+\sqrt x\right)} - \frac 5x\]

OpenStudy (paxpolaris):

rationalize denominator: \[\Large ={1\over 2 \left( x+ \sqrt x \right)}\cdot{\left( x- \sqrt x \right) \over\left( x- \sqrt x \right)} - \frac 5x\] \[\Large ={x-\sqrt x \over 2\left( x^2-x \right)}-\frac 5x\]

OpenStudy (paxpolaris):

\[\Large ={\left( x-\sqrt x \right)-10\left( x-1 \right) \over 2x(x-1)}\]

OpenStudy (anonymous):

The options given are \[10-9\sqrt{x}/2x(1+\sqrt{x})\] \[-10-9\sqrt{x}/2x(1+\sqrt{x})\] \[-10-9\sqrt{x}/2(1+\sqrt{x})\] \[-10-9\sqrt{x}/2x\]

OpenStudy (paxpolaris):

all right:\[\Large y\ ' ={1 \over 2 \sqrt x\left(1+\sqrt x\right)} - \frac 5x\]\[\Large ={\sqrt x \over 2 x\left(1+\sqrt x\right)} - {5\cdot2\left( 1+\sqrt x \right) \over 2 x\left(1+\sqrt x\right)}\]

OpenStudy (paxpolaris):

\[\huge = {-10-9\sqrt x \over 2x \left( 1+\sqrt x \right)}\]

OpenStudy (anonymous):

Thank you very much everyone. PaxPolaris special thanks!